The square of Sam's age increased by twice his age is 48. What is Sam's age?
I chose 2(a)2=48 and got a=4.9 years old. Am I right?

Answer by edjones(8007)

Let x be Sam's age.
x^2+2x=48
x^2+2x-48=0
(x+8)(x-6)=0
x=6 y.o.
.
Always check:
6^2+2*6=48
36+12=48 true
.
Ed